OpenStudy (anonymous):

Find a diameter: (15,15) and (-1,-1)

OpenStudy (anonymous):

47.14285714285714 radius = 1/2 of diameter radius=7.5 which is half of 15 using circumference of circle formula (2 * by pi * by radius) pi=22/7 2*22/7*7.5=47.14285714285714

OpenStudy (anonymous):

The diameter is the distance between these two points